# Bayesian Speech synthesizers Can Learn from Multiple Teachers

### v1

@@ Abstract
Codec-based text-to-speech (TTS) models have recently gained traction for their efficiency and strong performance in voice cloning. However, codec-based TTS faces limitations due to the challenges of pretraining robust speech codecs and the quality degradation introduced by quantization errors. Emerging evidence suggests that continuous-valued generative models can alleviate these issues and serve as a promising alternative. Yet, effectively modelling diverse speech patterns and developing reliable sampling strategies for continuous-valued autoregressive (AR) TTS remains underexplored. In this work, we propose BELLE, Bayesian evidential learning with language modelling for TTS, a novel continuous-valued AR framework that directly predicts mel-spectrograms from textual input. BELLE treats each mel-spectrogram frame as a Gaussian distribution sampled from a learned hyper-distribution, enabling principled uncertainty estimation, particularly in scenarios with parallel data (i.e., one text-audio prompt paired with multiple speech samples). To obtain such data, diverse speech samples are synthesized using multiple pre-trained TTS models given the same text-audio prompts, which are distilled into BELLE via Bayesian evidential learning. Experimental results indicate that BELLE demonstrates highly competitive performance compared with the current best open-source TTS models, even though BELLE is trained on a large amount of synthetic data and uses only approximately one-tenth of their training data. Audio samples generated by BELLE are available at https://belletts.github.io/Belle/. The code, checkpoints, and synthetic data will be released after the paper is accepted.
@@ 1 Introduction
Recently, autoregressive (AR) next-token prediction models based on discrete audio codecs have gained significant attention in the audio generation and text-to-speech (TTS) communities, driven by the success of models such as AudioLM Borsos et al. (2023) and VALL-E (Wang et al., 2023), among others. However, continuous-valued prediction has emerged as a promising alternative due to its ability to eliminate quantization errors introduced by token-based codecs (Meng et al., 2024; Liu et al., 2024; Lin & He, 2025; Chen et al., 2024b; Zhu et al., 2024; Eskimez et al., 2024; Chen et al., 2024a; Wang et al., 2025a; Jia et al., 2025; Lee et al., 2025). Generative models typically rely on sampling mechanisms to introduce stochasticity and improve synthesis quality. In audio-codec-based AR TTS, this is often achieved using mature top-$p$ sampling techniques, similar to those employed in token-based language models (LMs). Diffusion-based models, by contrast, introduce randomness by injecting noise into the input, thereby embedding variability directly into the generation process. However, effective sampling strategies for continuous-valued AR models remain underexplored. Recent studies have begun to address this gap. In image generation, MAR proposed a hybrid approach where an AR model predicts conditioning signals, followed by diffusion-based sampling (Li et al., 2024a). In zero-shot TTS, MELLE introduced Gaussian sampling (Meng et al., 2024), while FELLE adopted flow-matching techniques (Lipman et al., 2022; Wang et al., 2025a) to enhance sampling efficiency and output diversity.
In this paper, we propose BELLE: Bayesian evidential learning with language modelling for TTS Synthesis—a continuous-valued AR model that directly predicts mel-spectrograms from text inputs based on the evidential deep learning (EDL) framework (Amini et al., 2020). This represents the first application of Bayesian deep learning to TTS, enabling controllable and theoretically grounded sampling at test time. Specifically, BELLE models each predicted mel-spectrogram frame as a sample from a Gaussian distribution, whose parameters themselves are drawn from hyper-distributions. Metaphorically, the model first predicts the parameters of this hyper-distribution, then samples a specific Gaussian distribution from it, and finally generates mel-spectrogram frames from the sampled Gaussian. However, learning accurate Gaussian posterior estimates from limited data poses a challenge, as typical TTS datasets contain only a single audio recording per text prompt. To overcome this, the training corpus is augmented with synthetic audio samples generated by several publicly available pre-trained zero-shot TTS models (Du et al., 2024; Deng et al., 2025; Wang et al., 2025c; 2024; Casanova et al., 2024).
Our main contributions are as follows:
  • We present BELLE, the first AR TTS model using evidential deep learning to train and sample from continuous-valued Mel-Spectrogram space. The results show that Bayesian evidential sampling outperforms conventional Gaussian sampling.
  • We propose the first multi-teacher knowledge distillation approach for TTS model training, integrating it with evidential deep learning.
  • BELLE achieves competitive performance compared with leading open-source TTS systems trained on approximately 50k hours of speech data, while BELLE uses only around 5k hours of data that largely consist of synthetic samples. Additionally, its streaming variant, BELLE-stream, attains an effective balance between audio generation quality and latency.
@@ 2 Related Work
Continuous-valued TTS models can be broadly categorized into AR and non-autoregressive (NAR) approaches. AR methods include MELLE (Meng et al., 2024), which generates mel spectrogram frames using decoder-only Transformer-based AR modelling combined with Gaussian sampling. FELLE (Wang et al., 2025a), ARDiT (Liu et al., 2024) and DiTAR (Jia et al., 2025) first generate continuous-valued hidden states autoregressively, then employ flow-matching or diffusion-based modules, such as DiT (Peebles & Xie, 2023), to sample and produce mel spectrograms or a continuous latent space learned by a VAE (Pinheiro Cinelli et al., 2021). Moreover, VAE-based AR methods, including GMM-LM (Lin & He, 2025) and KALLE (Zhu et al., 2024), initially leverage a VAE to estimate a latent distribution (mean and variance), and subsequently autoregressively predict latent representations sampled from this distribution. NAR models such as E2 (Eskimez et al., 2024) and F5 (Chen et al., 2024a), on the other hand, predict mel-spectrograms directly via iterative mask-based refinement. Recently, several streaming TTS systems have emerged (Du et al., 2024; Yang et al., 2024; Sun et al., 2025; Sheng et al., 2025; Wang et al., 2025b). Most of these approaches model discrete features using a language model, followed by a flow-matching module or other similar components to generate speech features. Such architectures are relatively complex and may introduce potentially higher latency.
Bayesian Methods in Sequence Modelling and TTS.
Evidential Deep Learning (EDL) (Amini et al., 2020) is commonly used in recurrent neural networks (RNNs) and Transformers to provide per-time-step uncertainty (Cui et al., 2024; Khot et al., 2024; Marvi et al., 2025), benefiting tasks like time-series forecasting (Khot et al., 2024; Marvi et al., 2025). Other Bayesian approaches, such as Bayesian neural networks (Kononenko, 1989) in Long Short-Term Memory (LSTMs) and Transformers (Muñoz et al., 2024), and frameworks like SYMHnet (Abduallah et al., 2024), also address sequential uncertainty. Previous Bayesian TTS work focused on HMM-based models (Hashimoto et al., 2009b; a; Lu & King, 2012); to our knowledge, we are the first to apply Bayesian methods to neural network-based TTS.
Teacher-Student Training in TTS.
Knowledge distillation (KD) transfers knowledge from larger teacher models to smaller, faster student models (Nguyen et al., 2025; Yang et al., 2022; Li et al., 2024b). Originally, KD is used to reduce exposure bias in AR models like Tacotron (Wang et al., 2017; Liu et al., 2020) and to guide NAR models such as FastSpeech (Ren et al., 2019; 2020). It has been further applied to improve pronunciation (Nguyen et al., 2025), style transfer (Yang et al., 2022), perceptual quality in diffusion-based models (Li et al., 2024b), and to distill semantic knowledge from HuBERT (Gállego et al., 2025). To our knowledge, multi-teacher KD has not yet been explored in TTS.
@@ 3.2 General Architecture of Mel-based Autoregressive TTS Model
Figure 1: The structure of BELLE and the detailed sampling module. The output is assumed to follow a Normal-Inverse-Gamma (NIG) distribution, and the Sampling Module predicts four distribution parameters. Sequentially, variance and mean are obtained via Inverse-Gamma sampling and Gaussian sampling, respectively, followed by a final Gaussian sampling step to generate the output $\bm{y}_{t}^{(1)}$. BELLE-stream adopts the same architecture as BELLE. The text and Mel-spectrogram are split into chunks and interleaved to form the input sequence. Typically, the final chunk generates all remaining audio, making it longer than the preceding ones.
@@ 4.1 Preliminary: EDL in Mel-Spectrogram Prediction
$\bm{y}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}^{2}),\quad\bm{\mu}_{t}\sim\mathcal{N}\left(\bm{\gamma}_{t},\frac{\bm{\sigma}_{t}^{2}}{\bm{\nu}_{t}}\right),\quad\bm{\sigma}_{t}^{2}\sim\Gamma^{-1}(\bm{\alpha}_{t},\bm{\beta}_{t}),$ (3)
@@
where the NLL enforces distributional fit and the regularization term penalizes incorrect evidence. Details about EDL are in Appendix D.
@@ 4.2 From Gaussian Sampling to Bayesian Evidential Sampling
MELLE (Meng et al., 2024) directly predicts the mean and variance of a Gaussian distribution, then samples embeddings accordingly; implementation details are in Appendix E.1. We propose to replace the Gaussian sampling mechanism with a Bayesian evidential approach. Specifically, NIG distribution parameters are predicted from the AR hidden representation $\bm{e}_{t}$ via a linear layer followed by suitable activation functions to constrain each parameter within appropriate numerical ranges:
@@
$\bm{\sigma}_{t}^{2}\sim\Gamma^{-1}(\bm{\alpha}_{t},\bm{\beta}_{t}),\quad\bm{\mu}_{t}\sim\mathcal{N}\left(\bm{\gamma}_{t},\frac{\bm{\sigma}_{t}^{2}}{\bm{\nu}_{t}}\right),\quad\bm{z}_{t}\sim\mathcal{N}\left(\bm{\mu}_{t},\bm{\sigma}_{t}^{2}\right).$ (7)
@@ 4.3 Training Objectives
4.4 Multi Teacher Learning
@@
Let $\mathcal{L}(\bm{y}^{\text{gt}}_{i})$ denote the complete loss computation defined in Eqn. equation 8 using $\bm{y}^{\text{gt}}_{i}$ as the ground-truth mel-spectrogram. The final overall training loss is computed as a weighted sum:
$\mathcal{L}_{\text{multi-teacher}}=\sum\nolimits_{i=1}^{N+1}w_{i}\mathcal{L}(\bm{y}^{\text{gt}}_{i}),\quad\text{with}\quad\sum\nolimits_{i=1}^{N+1}w_{i}=1,\quad w_{i}\geq 0.$ (12)
@@ 5.1 Training Data and Details
The training dataset is derived from the Librispeech (Panayotov et al., 2015) training set and contains audio samples whose duration is from 0.5 second to 14 seconds, creating a final training dataset containing approximately 706 hours of speech. To provide richer acoustic diversity necessary for robust Bayesian distribution estimation, our training data are augmented by synthesizing multiple audio samples for each textual input using six publicly available pretrained TTS models, namely CosyVoice2 (Du et al., 2024), IndexTTS (Deng et al., 2025), SparkTTS (Wang et al., 2025c), F5TTS (Chen et al., 2024a), MaskGCT (Wang et al., 2024), and XTTS-v2 (Casanova et al., 2024). Including the TTS-synthesized data, the total dataset amounts to 4,817 hours of speech. Details of the data processing methodology can be found in Appendix C.1.
BELLE, BELLE-stream and MELLE are trained on the training dataset, following the training strategy of combining various data sources within each batch (see Sec. 4.4). For BELLE-stream, we set $S_{\mathrm{text}}=20$ and $S_{\mathrm{audio}}=50$, corresponding to an audio chunk duration of approximately $0.8$ seconds. It’s observed that the synthesis quality deteriorates with speech prompt. To address this, we first investigate the potential of BELLE for streaming synthesis in a single-speaker scenario, finetuning BELLE-stream using recordings from a single speaker of Librispeech to fix its voice timbre. Details about model configuration and training could be found in Appendix C.2 and C.3 respectively.

### v2

@@ Abstract
Text-to-Speech (TTS) is inherently a “one-to-many” mapping characterized by intrinsic uncertainty, yet current paradigms often oversimplify it into a deterministic regression task. While continuous-valued autoregressive (AR) models have recently emerged as a promising alternative to discrete codec-based approaches, they typically rely on a fixed-variance prior, fundamentally constraining generation to a static point estimate that ignores the dynamic variability of natural speech. To bridge this gap, we propose BELLE (Bayesian evidential learning with language modelling), a framework that shifts from deterministic prediction to principled Bayesian inference without increasing model parameters or inference latency. By modeling the acoustic target as a Normal-Inverse-Gamma distribution, BELLE captures data-dependent aleatoric uncertainty. To enable accurate variance estimation on standard single-reference datasets, we introduce a “one-to-many” training strategy that leverages synthetic samples as a statistical support set, allowing the model to learn robust distributional properties rather than merely imitating teacher artifacts. Experiments demonstrate that BELLE, trained on only $\sim$5k hours of data, outperforms leading open-source models trained on 50k hours (achieving a 25.8% relative WER reduction) and naturally supports high-quality streaming generation. Audio samples are available at https://belletts.github.io/Belle/.
TTS, EDL
@@ 1 Introduction
Text-to-Speech (TTS) synthesis is fundamentally a “one-to-many” mapping problem. In natural human speech, intrinsic uncertainty is ubiquitous: the same textual content can be spoken with substantial variations in prosody, rhythm, and acoustic details, even by the same speaker. Consequently, a robust TTS system should not merely predict a single deterministic output but rather model the conditional probability distribution of potential speech realizations. However, this stochastic nature is frequently overlooked by the research community, a blind spot reflected in both data availability and modeling paradigms. On the data side, standard TTS datasets predominantly adhere to a rigid “one-to-one” mapping—providing only a single acoustic recording per text—which effectively masks the intrinsic distributional variance of speech. Parallel to this, most existing models treat synthesis as a deterministic regression task or rely on heuristic sampling methods (e.g., top-$p$ or dropout) that lack a principled theoretical basis for modeling data-dependent uncertainty.
In parallel, the architecture of TTS models has evolved rapidly to improve generation fidelity. Autoregressive (AR) models based on discrete audio codecs have gained dominance, exemplified by AudioLM (Borsos et al., 2023) and VALL-E (Wang et al., 2023). While successful, these discrete approaches suffer from quantization errors and information loss inherent to tokenization. To address this, continuous-valued prediction has emerged as a promising alternative (Meng et al., 2024; Liu et al., 2024; Lin and He, 2025; Chen et al., 2024; Wang et al., 2025a), eliminating quantization artifacts by directly modeling continuous acoustic features (e.g., Mel-spectrograms). Despite the shift from discrete to continuous representations improves the representation of speech, it does not by itself address the missing probabilistic modeling of “one-to-many” variability, and probabilistic modeling in this domain remains stagnant. State-of-the-art models like MELLE (Meng et al., 2024) typically rely on a fixed-variance prior assumption. While effective for stability, this formulation fundamentally constrains the model to a point estimate, ignoring the complex, data-dependent aleatoric uncertainty inherent in expressive speech.
To bridge this gap, we propose BELLE (Bayesian evidential learning with language modelling) based on Evidential Deep Learning (EDL) (Amini et al., 2020). Unlike superficial architectural modifications, BELLE represents a substantive shift in modeling philosophy: moving from deterministic regression to principled Bayesian inference. Specifically, BELLE models each acoustic target with a Normal-Inverse-Gamma (NIG) distribution, dynamically predicting variance $\sigma^{2}$ conditioned on context. This capability enables the model to effectively capture the intrinsic “one-to-many” uncertainty of speech, facilitating grounded and diverse sampling that traditional regression objectives cannot achieve. Crucially, this Bayesian treatment integrates seamlessly into the architecture—improving robustness and naturalness without adding any parameters or computational overhead to the inference process compared to standard regression baselines.
Implementing this Bayesian framework, however, introduces a statistical challenge: accurately estimating the variance of a distribution theoretically requires multiple observations for a given input, whereas standard TTS datasets typically provide only a single recording per text. To resolve this, we augment the training set with synthetic samples from diverse pre-trained TTS systems. Crucially, these samples serve not as targets for imitation, but as a statistical support set to characterize the variance structure. By aggregating diverse realizations, our evidential framework learns robust distributional properties, effectively mitigating the influence of potential errors or artifacts from individual teacher models.
[…]
@@ 2 Related Work
Continuous-valued TTS models are primarily categorized into autoregressive (AR) and non-autoregressive (NAR) approaches. Among AR methods, MELLE (Meng et al., 2024) directly predicts mel spectrogram frames using a Transformer-based model with Gaussian sampling. Several recent studies have also explored continuous modeling by incorporating diffusion or flow-matching modules into AR frameworks (Wang et al., 2025a; Liu et al., 2024; Jia et al., 2025), or by leveraging VAE-based latent distributions (Lin and He, 2025; Zhu et al., 2024). In contrast, NAR models like E2 TTS (Eskimez et al., 2024) and F5-TTS (Chen et al., 2024) generate mel-spectrograms via iterative mask-based refinement. Regarding streaming TTS, recent systems (Du et al., 2024; Yang et al., 2024; Sun et al., 2025; Sheng et al., 2025; Wang et al., 2025b) typically combine discrete language models with flow-matching decoders. However, such hybrid architectures often introduce higher complexity and potential latency compared to purely autoregressive solutions.
Evidential Learning vs. Bayesian Neural Networks.
Bayesian modeling in deep learning is broadly divided into approaches that place priors over network weights—classic Bayesian Neural Networks (BNNs) (MacKay, 1992; Neal, 1996)—and those that place priors over the data likelihood parameters, known as Evidential Deep Learning (EDL) (Sensoy et al., 2018; Amini et al., 2020). Standard BNNs typically approximate the weight posterior via Variational Inference (Graves, 2011; Blundell et al., 2015) or Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016). While rigorous, these methods inherently require multiple forward passes or ensemble sampling to estimate predictive uncertainty, introducing significant latency that is prohibitive for real-time streaming TTS. In contrast, EDL adopts an Empirical Bayes (or Type-II Maximum Likelihood) framework. Rather than sampling weights, EDL treats the hyperparameters of the likelihood distribution (e.g., the Normal-Inverse-Gamma parameters) as the prediction targets, allowing the model to analytically marginalize over aleatoric uncertainty in a single deterministic forward pass. This distinction is critical for BELLE: it enables the rigorous modeling of the heteroscedastic, “one-to-many” nature of speech (as detailed in Appendix B) without the computational overhead of weight sampling, effectively balancing high-fidelity probabilistic modeling with the efficiency required for low-latency generation.
@@ 3.2 General Architecture of Mel-based Autoregressive TTS Model
Figure 1: The structure of BELLE and the detailed sampling module. The output is assumed to follow a Normal-Inverse-Gamma (NIG) distribution, and the Sampling Module predicts four distribution parameters. Sequentially, variance and mean are obtained via Inverse-Gamma sampling and Gaussian sampling, respectively, followed by a final Gaussian sampling step to generate the output $\bm{y}_{t}^{(1)}$.
@@ 3.2.4 Streaming Chunk-based Generation
Figure 2: BELLE-stream adopts the same architecture as BELLE. The text and Mel-spectrogram are split into chunks and interleaved to form the input sequence. Typically, the final chunk generates all remaining audio, making it longer than the preceding ones.
@@ 4.1 Preliminary: EDL in Mel-Spectrogram Prediction
$\begin{gathered}\bm{y}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}^{2}),\quad\bm{\mu}_{t}\sim\mathcal{N}\left(\bm{\gamma}_{t},\frac{\bm{\sigma}_{t}^{2}}{\bm{\nu}_{t}}\right)\\ \bm{\sigma}_{t}^{2}\sim\Gamma^{-1}(\bm{\alpha}_{t},\bm{\beta}_{t}),\end{gathered}$ (3)
@@
where the NLL enforces distributional fit and the regularization term penalizes incorrect evidence. Details about EDL are in Appendix E.
@@ 4.2 From Gaussian Sampling to Bayesian Evidential Sampling
In the MELLE framework (Meng et al., 2024), the latent embedding generation is modeled as a stochastic process. At each timestep $t$, the model predicts the parameters of a multivariate Gaussian distribution from the autoregressive hidden representation $\bm{e}_{t}$. Specifically, a linear projection is used to regress the mean $\bm{\mu}_{t}$ and the diagonal log-variance $\log\bm{\sigma}_{t}^{2}$:
To enable differentiability during training, the latent embedding $\bm{z}_{t}$ is synthesized using the reparameterization trick, expressed as $\bm{z}_{t}=\bm{\mu}_{t}+\bm{\sigma}_{t}\odot\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$.
A critical component of MELLE is its training objective, which employs a Kullback-Leibler (KL) divergence loss to align the predicted distribution $p_{\bm{\theta}}(\bm{z}_{t}|\bm{e}_{t})$ with a target prior distribution $p(\bm{z}_{t})$. MELLE defines the prior distribution centered at the ground-truth acoustic feature $\bm{y}_{t}$ with a fixed unit variance:
$p(\bm{z}_{t})=\mathcal{N}(\bm{y}_{t},\bm{I}).$ (7)
Consequently, the sampling loss is formulated as:
$\mathcal{L}_{\text{KL}}$ $=\sum_{t=0}^{T-1}D_{\text{KL}}\left(\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}^{2}\bm{I})\parallel\mathcal{N}(\bm{y}_{t},\bm{I})\right)$
$=\frac{1}{2}\sum_{t=0}^{T-1}\left(\|\bm{\mu}_{t}-\bm{y}_{t}\|^{2}_{2}+\|\bm{\sigma}_{t}\|^{2}_{2}-\sum_{i=1}^{D}\log\bm{\sigma}_{t,i}^{2}-D\right),$ (8)
where $D$ denotes the feature dimensionality. This formulation effectively serves a dual purpose: the term $\|\bm{\mu}_{t}-\bm{y}_{t}\|^{2}_{2}$ acts as a regression loss forcing the predicted mean to track the ground truth, while the variance terms regularize the model uncertainty towards unity, preventing the variance from collapsing to zero.
We propose to replace the Gaussian sampling mechanism with a Bayesian evidential approach. Specifically, NIG distribution parameters are predicted from the AR hidden representation $\bm{e}_{t}$ via a linear layer followed by suitable activation functions to constrain each parameter within appropriate numerical ranges:
@@
$\begin{gathered}\bm{\sigma}_{t}^{2}\sim\Gamma^{-1}(\bm{\alpha}_{t},\bm{\beta}_{t}),\quad\bm{\mu}_{t}\sim\mathcal{N}\left(\bm{\gamma}_{t},\frac{\bm{\sigma}_{t}^{2}}{\bm{\nu}_{t}}\right),\\ \bm{z}_{t}\sim\mathcal{N}\left(\bm{\mu}_{t},\bm{\sigma}_{t}^{2}\right).\end{gathered}$ (10)
Theoretical Comparison
The transition to Bayesian evidential learning offers fundamental theoretical advantages over the standard Gaussian objective. While MELLE minimizes KL divergence against a fixed-variance prior ($\mathcal{N}(\bm{y}_{t},\bm{I})$), implicitly assuming homoscedasticity (uniform uncertainty), our framework adopts an Empirical Bayes approach to perform Type-II Maximum Likelihood Estimation by placing a conjugate NIG prior over the likelihood parameters. This allows the model to dynamically estimate variance from evidence, effectively capturing the inherent heteroscedasticity of speech without increasing backbone parameters. Furthermore, marginalizing over this prior yields a Student-$t$ predictive distribution which strictly subsumes the Gaussian as a limiting case, thereby guaranteeing superior expressive capacity while maintaining analytically tractable, closed-form sampling (refer to Appendix B for details).
@@ 4.3 Training Objectives
4.4 “One-to-Many” Training Strategy
@@
Let $\mathcal{L}(\bm{y}^{\text{gt}}_{i})$ denote the complete loss computation defined in Eqn. (11) using $\bm{y}^{\text{gt}}_{i}$ as the ground-truth mel-spectrogram. The final overall training loss is computed as a weighted sum:
$\begin{gathered}\mathcal{L}_{\text{one-to-many}}=\sum\nolimits_{i=1}^{N+1}w_{i}\mathcal{L}(\bm{y}^{\text{gt}}_{i}),\\ \quad\text{with}\quad\sum\nolimits_{i=1}^{N+1}w_{i}=1,\quad w_{i}\geq 0.\end{gathered}$ (15)
@@ 5.1 Training Data and Details
The training dataset is derived from the Librispeech (Panayotov et al., 2015) training set and contains audio samples whose duration is from 0.5 second to 14 seconds, creating a final training dataset containing approximately 706 hours of speech. To provide richer acoustic diversity necessary for robust Bayesian distribution estimation, our training data are augmented by synthesizing multiple audio samples for each textual input using six publicly available pretrained TTS models, namely CosyVoice2 (Du et al., 2024), IndexTTS (Deng et al., 2025), SparkTTS (Wang et al., 2025c), F5TTS (Chen et al., 2024), MaskGCT (Wang et al., 2024), and XTTS-v2 (Casanova et al., 2024). Including the TTS-synthesized data, the total dataset amounts to 4,817 hours of speech. Details of the data processing methodology can be found in Appendix D.1.
BELLE, BELLE-stream and MELLE are trained on the training dataset, following the training strategy of combining various data sources within each batch (see Sec. 4.4). For BELLE-stream, we set $S_{\mathrm{text}}=20$ and $S_{\mathrm{audio}}=50$, corresponding to an audio chunk duration of approximately $0.8$ seconds. Details about model configuration and training could be found in Appendix D.2 and D.3 respectively.